Find the area of the region that is bounded by the given curve and lies in the specified sector. p=el-H), 5051
Answer #1

Solve: Gilveni rad) & II Lost . The area under Carne :- A= Sokolda Isolate o for: - If f(x) = g(x) then In (f(x)) = in /a(x))so; Suvl=ux-sur- Here, u=lult) & 4/21 so; oltre (mer) u~ t . dluerlahat also, wat adual dvaldr Sax=slar [ Jadnaartch Nagy SoRefine! - - - 21+ 4114 (t)- 2rln (1)+c Therefore, Finally, The area isi- A = – 2714 471n(t)-2114 (2) - 5.26 453... x 5.26FIGURE: - ?2077